You are looking into an investment that will pay you ZMW 12, 000 per year for the next 10 years. If you require a 15 percent at a return, what is the most you would pay for this investment?
1
Expert's answer
2020-10-20T05:23:48-0400
"Solution"

"PVIAF=\frac{1-(1+r)^{-n}}{r}\\ 
Where, PVIAF\ is\ the\ Present\ Value\ Interest\ Annuity\ Factor.\\
r=15\%, \ n=10\\
Therefore,\\
PVIAF=\frac{1-(1+0.15)^{-10}}{0.15}\\
=5.02\\
But, Present\ Value=PVIAF\times\ Future\ Value\\
PV=5.02\times12,000\\
=ZMW\ 60,240"